OpenStudy (paulalovee):

solve....-5/8c=20 ill medal n fan... -34, 34 -32 32

OpenStudy (jordanloveangel):

(-5/8) c = 20 - 5 c = 160 5 c = - 160 c = - 32

OpenStudy (paulalovee):

thx so much lol

OpenStudy (jordanloveangel):

Np gul